1) Let,

M = Event an officer is a man

W = Event an officer is a woman

A = Event an officer is promoted

B = Event an officer is not promoted

Suppose, a workforce consists of 1200 workers: 940 men and 240 women. Over the past two years, 324 officers were promoted: 288 males and 36 females.

  1. What is the probability that a randomly selected officer is a man and is promoted?
  2. What is the probability that a randomly selected officer is a man and is not promoted?
  3. What is the probability that a randomly selected officer is a woman and is promoted?
1
Expert's answer
2021-05-26T03:41:42-0400

1.)  The probability that a randomly selected officer is a man and is promoted "= \dfrac{288}{1200}"


2.) The probability that a randomly selected officer is a man and is not promoted "= 1- \dfrac{288}{1200}"


3.) The probability that a randomly selected officer is a woman and is promoted "= \dfrac{36}{1200}"